### v2

# LEGO Co-builder: Exploring Fine-Grained Vision-Language Modeling for Multimodal LEGO Assembly Assistants

Abstract
Vision-language models (VLMs) are facing the challenges of understanding and following multimodal assembly instructions, particularly when fine-grained spatial reasoning and precise object state detection are required. In this work, we explore LEGO Co-builder, a hybrid benchmark combining real-world LEGO assembly logic with programmatically generated multimodal scenes. The dataset captures stepwise visual states and procedural instructions, allowing controlled evaluation of instruction-following, object detection, and state detection. We introduce a unified framework and assess leading VLMs such as GPT-4o, Gemini, and Qwen-VL, under zero-shot and fine-tuned settings. Our results reveal that even advanced models like GPT-4o struggle with fine-grained assembly tasks, with a maximum F1 score of just 40.54% on state detection, highlighting gaps in fine-grained visual understanding. We release the benchmark, codebase, and generation pipeline to support future research on multimodal assembly assistants grounded in real-world workflows.
Introduction
Multimodal instruction-following assistants are gaining increasing relevance in domains requiring precise procedural understanding, such as furniture construction (You et al. 2022), automotive manufacturing (Bellalouna et al. 2020), and industrial product assembly (Funk et al. 2017). These tasks demand step-by-step reasoning, spatial awareness, and accurate interpretation of visual and textual instructions—capabilities that current AI systems still struggle to reliably deliver.
Figure 1: An example of incorrect state detection by GPT-4o during LEGO assembly. Given an image and text instructions, the model fails to accurately recognize an incorrectly placed LEGO part (highlighted in green), demonstrating the challenge of fine-grained vision-language alignment.
Figure 2: An illustration of the proposed three VLM tasks, highlighting the core capabilities required by the one-for-all VLM architecture for LEGO assembly based on the procedural instruction manual.
Traditional computer vision research has tackled various purely visual tasks, including action segmentation, recognition, anticipation, and object state detection (Damen et al. 2018; Miech et al. 2020; Wang et al. 2023b; Gao et al. 2024). Procedural activity understanding has also been explored using video datasets with or without textual narration (Tang et al. 2019; Zhukov et al. 2019; Li et al. 2023a; Sener et al. 2022). Some methods incorporate textual alignment, such as narration-video grounding or instruction-conditioned generation (Padmakumar et al. 2021; Miech et al. 2019), but often rely on separate architectures for each task type.
[…]
VLM offer a promising solution by bridging visual and textual modalities, enabling models to align procedural language with visual observations. Recent advances in large-scale models such as GPT-4o (Achiam et al. 2023), LLaVa (Liu et al. 2023), Qwen-VL (Bai et al. 2023), BLIP-2 (Li et al. 2023c), MiniGPT-v2 (Chen et al. 2023), and more recent reasoning-augmented variants like VLM-R1 (Shen et al. 2025), GLM-4.1v-thinking (Hong et al. 2025), and Gemini2.5 (Comanici et al. 2025), have demonstrated strong performance on a range of general-purpose multimodal tasks. However, most existing benchmarks focus on coarse-grained capabilities such as object recognition, captioning, and visual question answering, and do not evaluate the detailed, procedural understanding required for step-by-step manual assembly. While recent VLMs research (Laurençon et al. 2024; Cheng et al. 2024; Chen et al. 2024a) aim to unify vision-language reasoning under a single architecture, challenges remain—particularly for tasks that require dense and fine-grained perception (Wei et al. 2024; Rahmanzadehgervi et al. 2024; Chen et al. 2024b).
In this work, we address the need for instruction-following benchmarks that evaluate a model’s ability to reason over both visual and linguistic input in detail. We focus on LEGO brick assembly as a representative manual task that is richly structured, visually complex, and procedurally grounded. Figure 1 illustrates the limitations of current models in such tasks. When prompted with a visual scene and corresponding instruction, GPT-4o fails to detect a misplaced LEGO piece, highlighting the challenge of fine-grained visual scene understanding and object state recognition. This motivates the need for benchmarks that go beyond high-level visual understanding and evaluate precise spatial compliance with procedural steps.
To address this gap, we present LEGO Co-builder, a benchmark and dataset designed to evaluate fine-grained instruction-following capabilities in LEGO assembly tasks. Our dataset includes multimodal sequences consisting of visual snapshots, object states, and procedural steps derived from human-crafted manuals. We introduce a unified task formulation and evaluate nine leading VLMs in both zero-shot and fine-tuned settings. Our contributions are summarized as follows: (1) We develop a unified vision-language architecture to benchmark fine-grained, multimodal, instruction-following capabilities in procedural manual-guided assembly tasks. (2) We evaluate prevailing VLMs on a dataset of LEGO assembly sequences with grounded visual and textual supervision, under both zero-shot and fine-tuned settings. (3) We release the dataset, benchmarking code, and a modular synthetic data generation pipeline to support future research in multimodal instruction-following.
Societal Impact. This work advances the development of multimodal AI assistants that can transform how people learn and perform complex physical tasks. By enabling fine-grained vision-language understanding, our benchmark supports the creation of multimodal educational tools that combine visual input, language, and potentially augmented reality (AR) to enhance hands-on learning experiences. These systems can improve training in domains like education and industrial training. Moreover, the ability to interpret and verify procedural steps visually opens up promising avenues for assistive technologies—particularly for blind or visually impaired learners—by providing real-time, AI-driven guidance through tasks that were previously inaccessible. In this way, our research supports both innovation in teaching modalities and greater inclusivity in skill development.
Fine-Grained Vision-Language Modeling
Task Definition
We investigate manual-guided LEGO assembly and define the following three vision-language tasks (See in Figure 2):
  1. (T1)
    Scene Understanding. Given the current step’s image $V_{t}$ and the task-specific query $Q_{t}$, the model outputs a response containing the scene description for the current assembly step $I_{t}$. $Q_{t}$ is the task description contextualized with the previous step’s textual instruction $I_{t-1}$. It assesses the model’s ability to accurately describe the current assembly step from an instruction manual, including but not limited to generating object representations, their properties, and the assembly procedure.
  2. (T2)
    Object Detection. Given a seeking step’s image $V_{t^{\prime}}$ ($t^{\prime}=t-1$) and task-specific query $Q_{t^{\prime}}$, the model outputs a response containing the positional text $P_{t^{\prime}}$, which includes identified object and its coordinate formatted as “[Object] [Xleft] [Ytop] [Xright] [Ybottom]”. The query $Q_{t^{\prime}}$ is the task description contextualized with the corresponding textual instruction $I_{t^{\prime}}$. It assesses the model’s ability to accurately identify and locate objects in the scene, ensuring they are positioned correctly for the next assembly action.
  3. (T3)
    State Detection. Given an assembly step’s image, either the original manual image $V_{t}$ or a generated negative sample $\tilde{V}_{t}$, and task-specific query $Q_{t}$, the model outputs a response indicating a correct or incorrect state. It evaluates the model’s ability to verify the accuracy of the assembly progression, determining if the assembly action has been correctly completed.
These tasks have been crafted to rigorously test and demonstrate the capabilities of VLMs in LEGO Assembly, focusing on their ability to understand and interpret complex multimodal inputs.
One-for-all Vision-Language Architecture
We investigated existing VLMs and derived a universal architecture for the proposed tasks, as illustrated in Figure 3. Given an image, either from manuals or real-world scenes, and a task-specific query, a VLM based on this architecture generates a textual response as the output. This architecture integrates a vision encoder, a large language model, and a vision-language projector as core modules, along with a task-specific query for task adaptation. Specifically, the vision encoder processes visual inputs, the large language model interprets and generates textual information, and the vision-language projector aligns visual and textual data for seamless task execution. The query can be customized using the following key components: (1) Task-specific token:Special tokens such as “[grounding]”, “[object]”, and “[state]” are introduced for task T1, T2, T3, respectively, to enhance task focus and accuracy. (2) Instruction:Relevant manual instructions or task directives are incorporated to provide the model with context, ensuring responses align with task requirements. (3) Format-directive:A directive to clarify the expected output format, ensuring outputs are precise and directly applicable.
Figure 3: The architecture for one-for-all vision-language modeling. It integrates a vision encoder, a large language model, and a vision-language projector as core functional modules, with a task-specific query for task adaptation.
Dataset Creation
Manual Crawling and Scene Data Matching
We outline the procedure for creating the dataset for task T1. First, we collected 65 official LEGO instruction manuals (LEGO 2024), designed to help blind and visually impaired users assemble LEGO sets accurately. Each instruction manual comprises several elements: (1) step-by-step textual instructions; (2) corresponding image for each step; and (3) tags provided by LEGO include instruction and image tags. Then, we split a full manual into several assembly sessions, each session includes two types of steps: object seeking and object assembly. Each seeking step is followed by an assembly step , forming a pair. Next, we iterated through all sessions to generate the scene understanding dataset, $D_{T1}=\{(Q_{t},V_{t},A_{t})\}|_{t=0}^{|D_{T1}|}$, where each element is a triplet of (query, image, text), and the query $Q_{t}$ is constructed by filling the query template with the task token “[grounding]” and the previous instruction $I_{t-1}$ as context.
Object Position Inference
This subsection outlines the procedure for creating the dataset for task T2. First, we iterated through all object-seeking steps, where users are asked to find specific objects. Second, we iterated all images to generate the object detection dataset, $D_{T2}=\{(Q_{t^{\prime}},\tilde{V}_{t^{\prime}},A_{t^{\prime}})\}|_{t^{\prime}=0}^{|D_{T2}|}$, where each element is a triplet of (query, image, text) for a seeking step $t^{\prime}=t-1$. The query $Q_{t^{\prime}}$ is constructed by filling the query template with the task token “[object]” and the current instruction $I_{t^{\prime}}$. The response $A_{t^{\prime}}$ contains a positional text $\tilde{P}_{t^{\prime}}$ formatted as “[Object] [Xleft] [Ytop] [Xright] [Ybottom]”. This is initially generated by querying the image using MiniGPT-v2 (Chen et al. 2023), the state-of-the-art model at the time of this work. The composite image $\tilde{V}_{t^{\prime}}$ is created by combining the current image ${V}_{t^{\prime}}$ with three randomly sampled images. Last, the initial coordinates are adjusted to fit the composite image.
Variant State Generation
This subsection outlines the procedure for creating the dataset for task T3. First, we iterated through all object-assembly steps, along with their corresponding previous steps, where users are trained to assemble the parts identified in the prior steps. Second, we conducted part segmentation by detecting the boundary color and segmenting the part to be assembled. In each assembly session, the objects to be assembled in the current step are tagged in the previous step, with their boundaries highlighted in distinct colors that vary between sessions. The highlighted colors are selected based on predominant color detection using K-means clustering (Statsmodels 2024), followed by Hue filtering (Chu, Tsuji, and Kato 2014) and manual correction for accuracy. Third, we added bounding boxes to the segmented objects. Fourth, we applied a natural perturbation to the part to be assembled by randomly shifting it within the background box, ensuring it moves by at least 5%. We treated each state from the manual as a positive sample and generated three variant states as negative samples. Then, we constructed the state detection dataset, $D_{T3}=\{(Q_{t},V_{t},A_{t},\tilde{V}_{t},\tilde{A}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}},\tilde{A}_{t}^{{}^{\prime\prime}})\}|_{t=0}^{|D_{T3}|}$, where the query $Q_{t}$ is constructed by filling the query template with the task token “[state]”, the current instruction $I_{t}$, and the previous instruction $I_{t-1}$ as context. The variant states $\{\tilde{V}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}}\}$ are natural pertubations of $V_{t}$. The responses $A_{t}$ and $\{\tilde{A}_{t},\tilde{A}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime\prime}}\}$ indicate the correct and incorrect statuses of the assembly state, respectively.
Dataset Statistics
We summarize the statistics of LEGO-VLM dataset in Table 1. It is generated from 65 LEGO instruction manuals and divided into 397 sessions, covering 5,612 scenes, 4,784 objects, and 2,716 states. It consists of 5,614 instruction steps, with each step containing an image and corresponding textual instructions. Out of these, 4,814 steps focus on object seeking, while 5,614 steps involve parts assembly, covering 3,172 states with 222 distinct boundary colors. Overall, the dataset includes 35,612 vision-language data samples: 5,612 for T1, 19,136 for T2, and 10,864 for T3, respectively.
# Manual # Session # Scene # Object # State
65 397 5,612 4,784 2,716
# Step Overall Identification Assembly
10,428 4,814 5,614
# Sample Overall T1 T2 T3
35,612 5,612 19,136 10,864
Table 1: Statistics of LEGO-VLM dataset.
Experimental Setup
Benchmarks
We benchmark the following nine prevailing VLMs: (1) mPLUG-OWL2 (Ye et al. 2024)replaces attention with a modality adapter in a large language model (LLM) decoder. (2) BLIP2 (Li et al. 2023c)uses a two-stage pre-trained Q-Former between an image encoder and a LLM to bridge the vision-language modality gap. (3) LLaVa (Liu et al. 2023)combines a visual CLIP encoder and a language decoder Vicuna for general-purpose visual language understanding. (4) Qwen-VL (Bai et al. 2023)connects a LLM with a visual encoder using position-aware vision-language adapter towards fine-grained visual understanding. (5) InstructBLIP (Dai et al. 2023)explores general-purpose vision-language instruction tuning based on the pretrained BLIP2. (6) MiniGPT-v2 (Chen et al. 2023)links a frozen ViT visual encoder with Llama-2 via a projection layer, applicable for diverse tasks via task-specific multimodal instructions. (7) Otter (Li et al. 2023a)is tuned on the OpenFlamingo, conditioning the language model on images for multi-modal perception and reasoning. (8) MiniGPT-4 (Zhu et al. 2023)integrates a frozen visual ViT&Q-Former encoder and Vicuna via a projection layer, unlocking advanced multimodal capabilities like GPT-4. (9) GPT-4o(OpenAI 2024)is a widely used commercial model accessible via APIs. (10) Qwen-VL-2.5update Qwen-VL with the latest Qwen LLM (QwenTeam 2024). (11) Gemini-2.5-flash(Comanici et al. 2025)is a thinking model, designed to tackle increasingly complex problems. (12) GLM-4.1-thinking(Hong et al. 2025)is designed to explore the upper limits of reasoning.
Evaluation Metrics
[…]
Outcomes
Benchmark Results
We compare the performance of nine prevailing VLMs on the proposed LEGO-VLM dataset, without and with finetuing, as shown in Table 2.
Model T1: Scene understanding T2: Object detection T3: State detection
F1-Theme $\uparrow$ ROUGE $\uparrow$ BLEU $\uparrow$ F1-Object $\uparrow$ IOU $\uparrow$ F1-State $\uparrow$ FPR $\downarrow$
PEFT (LoRA) w/o w/ w/o w/ w/o w/ w/o w/ w/o w/ w/o w/ w/o w/
mPLUG-OWL2 23.16 29.70 25.23 32.75 3.04 8.46 77.94 93.05 14.23 34.88 36.82 15.00 63.41 35.00
BLIP2 27.85 32.65 29.55 40.35 6.40 12.50 77.35 84.48 34.25 40.57 24.71 28.50 100.00 94.87
LlaVA 32.10 35.88 13.63 42.97 0.99 17.73 54.39 72.32 $\emptyset$ 60.98- 25.04 30.00 99.41 50.00
Qwen-VL 34.84 37.00 29.57 39.08 5.19 13.13 78.56 89.15 25.60 30.08 39.77 39.53 97.99 96.82
InstructBLIP 32.85 32.85- 29.20 36.87 4.91 11.52 79.76 98.16 $\emptyset$ 47.20- 0.00 0.02 1.22 0.00
MiniGPT-v2 33.06 37.52 34.72 32.56 9.81 8.28 84.95 85.91 26.98 25.94 36.76 38.64 60.42 80.72
Otter 11.49 / 12.12 / 1.28 / 72.39 / $\emptyset$ / 35.33 / 75.88 /
MiniGPT-4 34.11 / 15.05 / 1.88 / 87.09 / 30.20 / 37.19 / 67.37 /
GPT-4o 25.81 / 18.67 / 2.00 / 66.67 / 21.68 / 40.54 / 43.06 /
Gemini-2.5-flash 2.31 / 28.28 / 6.21 / 93.12 / 13.16 / 28.81 / 8.72 /
Qwen-VL-2.5 19.13 / 15.34 / 2.70 / 75.55 / 10.63 / 33.85 / 36.70 /
GLM-4.1-thinking 35.73 / 19.44 / 2.31 / 70.64 / 4.40 / 39.61 / 38.53 /
Table 2: Benchmarking vision-language models on LEGO-VLM dataset. The upper part represents open-resourced vision-language models, both without (/wo) and with (/w) PEFT using LoRA. The lower part presents results obtained via API calls to the latest vision-language models. The bold font indicates the highest score in each column. Symbols $\uparrow$ and $\downarrow$ denote that higher and lower values are better, respectively. Symbol “-” indicates the model is not applicable for fine-tuning. Symbol $\emptyset$” denotes a meaningless zero as the model fails to generate output as instructed. The superscripts “”, “”, and “-” indicate an increase, decrease, or inapplicability in the evaluation score after fine-tuning, respectively.
First of all, existing models struggle with fine-grained assembly tasks in AR. While the F1-Object score for task T2 using the fine-tuned InstructBLIP is high at 98.16%, the intersection over union (IOU) is only 47.20%, indicating minimal overlap between the predicted object positions and the ground truth, which is still unsatisfactory. Even the commercial model GPT-4o achieves only 66.67% for the F1-Object score and 21.68% on task T2. This might also caused by the difficulty of understanding scenes. For example, fine-tuned LlaVA shows the best overlap with reference instructions, scoring 42.97% on ROUGE and 17.73% on BLEU. However, the fine-tuned MiniGPT-v2, despite being the top performer in theme entity identification, achieves only 37.52%. This indicates that understanding and generating theme entities, such as LEGO parts and their properties, remains a significant challenge. For state detection in task T3, the fine-tuned InstructBLIP achieves a 0.00% of FPR, meaning it almost perfectly avoided incorrectly classifying negative instances as positive. However, the F1-State score is just 0.02%, indicating a significant failure in identifying the correct states. These results highlight the challenges of the proposed fine-grained vision tasks, reinforcing the need to advance this area and its resources as a valuable research topic.
Second, open-resource VLMs with fine-tuning generally outperform or are comparable to commercial models. For one hand, top-performing open-resource VLMs are more informative with identified accurate entities. For example, MiniGPT-v2 with fine-tuning achieves an 11.71% higher F1-Theme score on task T1; InstructBLIP with fine-tuning achieves a 31.48% higher F1-Object score, while Qwen-VL shows only a 1% decrease, compared with the commercial GPT-4o. This may be due to the proposed dataset enhancing the model’s understanding of domain-specific knowledge, such as LEGO parts and their properties. On the other hand, top-performing open-resource VLMs can generate assembly instructions that closely align with the provided manuals. For example, the fine-tuned LlaVA achieves ROUGE and BLEU scores that are 2.30 and 8.87 times higher than those of GPT-4o, respectively. Besides, fine-tuning generally improves the performance of VLMs, as indicated by the results marked with the superscript “”. This highlights the significant difference between the proposed dataset and those used to train general commercial models.
Last, task difficulty varies significantly, with the biggest challenge being the alignment of positional information in images with the corresponding textual information from the query. Specifically, in comparison to task T1, tasks T2 and T3 present greater challenges in instruction following. For example, several VLMs (i.e., LlaVA, InstructBLIP, Otter), without fine-tuning, fail to follow instructions to generate positional information for evaluating IOU, indicated by the symbol $\emptyset$.” Additionally, all evaluated VLMs have F1-State scores below 50%, indicating that their predictions are even worse than random guessing. One potential reason task T3 is complex, and relies on task T2 is that it requires positional information to detect states, as well as insights from task T1 to understand the LEGO components and their relationships in the scene. We will explore this in future work, as this study primarily focuses on proposing the tasks rather than complex modeling.
Data Quality Assessment
To ensure the quality of the generated data for tasks T2 and T3, we sampled 100 data samples and conducted quality assessments for each task. We added bounding boxes based on the coordinates and asked three annotators to evaluate the generated data quality, focusing on coordinates, entities, and negative samples, based on the following criteria: (1) Entity disambiguitymeasures how clearly a target entity is relevant to the scene in T2, with scores of 0, 1, and 2 indicating low, medium, and high disambiguation, respectively. (2) Boundary precisionmeasures how accurately a bounding box encloses a target object in T2, with scores of 0, 1, and 2 indicating low, medium, and high precision, respectively. (3) State relevancemeasures whether the generated parts in an image are a relevant variation of the original image in T3. (4) State identifiabilitymeasures whether the generated parts in an image are a recognizable variation of the original image in T3.
Figure 4: Human assessment of data quality.
As shown in Figure 4, the average entity disambiguation and boundary precision scores are 1.74 and 1.25 out of 2 with the Kappa values (Scikit-learn 2024) of 0.60 and 0.61, respectively.
Query: [grounding] After I find 1 transparent mailbox front 2x2. Tell me what to do for the current situation. [Uncaptioned image]
Reference of Assembly Instruction:
Insert the two tabs at the bottom of the mailbox casing into the two small notches on the mailbox front.
mPLUG-OWL2: Place the mailbox front 2x2 next to the block on the table   .
BLIP2: Put the mailbox front 2x2 horizontally on the table, clasp to the back. Put the mailbox back 2x2 vertically on top, slope to the left.
LlaVA: In the image, there is a small Lego-like object with a red liquid inside of it. The object is placed in a blue sky, and it appears to be a 3D model. To create a transparent mailbox, you need to find a transparent mailbox front that is 2x2 in size. Once you find it, you can use it as the base for your transparent mailbox. You can then place the small Lego-like object inside the mailbox to complete the scene.
Qwen-VL: Put the transparent mailbox front 2x2 on the right side of the white arch.
InstructBLIP: Place it horizontally on the table, clasp to the back.
MiniGPT-v2: Here’s what you need to do with the pieces you have collected: 1. Place the transparent mailbox front 2x2 on top of the plate 1x2, aligned with the short side of the plate. Now you have completed the current situation. Let me know if you need help with anything else!
Otter: In view of the reasoning ‘A mailbox is small and square with a see through window.’, the optimal answer should be: build structure.
MiniGPT-4: Attach the transparent mailbox front to the base structure, aligning the tabs with the notches.
GPT-4o: Ensure the transparent mailbox front 2x2 is securely attached by inserting the small shaft into the round brick 1x1 before proceeding with the next step.
Table 3: An example for understanding VLM challenges.
This suggests that annotators reached a moderate agreement on the clarity of the evaluated objects in relation to the corresponding scenes. The average state relevance and identifiability scores are 0.93 and 0.74 out of 1.00, respectively, with the Kappa scores 0.81 and 0.59. This indicates that annotators reached near-perfect agreement on the relevance of the generated states to the original while showing moderate agreement on the recognizable variation of those states. Besides, the first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria. This ensures the overall quality of the generated states.
Case Study of Fine-grained VLM Challenges
In this section, we present illustrative examples to highlight the capabilities and challenges in fine-grained vision-language understanding, particularly in achieving overlap with the reference and accurately recognizing entities and their properties, such as size, color, and etc.
Table 3 present an example showcasing the intuitive results of the prevailing VLMs on scene understanding. Compared with the reference, existing VLM models face key challenges: (1) Vague and generic assembly instructions.mPLUG-OWL2 and BLIP2 produce broad, non-specific placement directives (e.g., “Place the mailbox front 2x2 next to the block on the table” or “Put the mailbox front 2x2 horizontally on the table, clasp to the back”), falling short of the precise, step-by-step guidance found in the reference. Otter’s response is even more superficial, merely stating “build structure” without actionable detail. (2) Struggles with fine-grained entity recognition and attribute grounding.While models like Qwen-VL and MiniGPT-v2 correctly mention the “transparent mailbox front 2x2,” others, such as InstructBLIP, overlook critical attributes (e.g., referring only to “it” or “the mailbox front”). LlaVA, despite its verbosity, hallucinates details (e.g., “red liquid inside” and “blue sky”) absent from the scene, highlighting confusion in entity identification and property association. (3) Tendency to hallucinate or misinterpret assembly steps.MiniGPT-v2 and GPT-4o introduce extraneous or incorrect actions, such as “aligning with the short side of the plate” or “inserting the small shaft into the round brick 1x1,” which are not part of the actual assembly process. This reveals a shallow grasp of both scene context and the logical flow of assembly operations. Overall, while some models can partially identify the correct entities and occasionally mention relevant properties, none consistently generate instructions that are both factually accurate and contextually appropriate for fine-grained assembly tasks.
Related Work
Vision-Language Datasets for Assembly Tasks
Various datasets have been developed to enhance assembly tasks by supporting key capabilities T1, T2, and T3. A comparative summary is presented in Table 4.
      Dataset       T1       T2       T3       Size
      COIN       ✓       ✗       ✓       11,827
      HowTo100M       ✓       ✗       ✓       23,611
      TEACh       ✓       ✗       ✓       3,215
      MIMIC-IT       ✓       ✗       ✗       2.8M
      HoloAssist       ✓       ✗       ✓       350
      EPIC-KITCHENS       ✗       ✓       ✓       89,977
      Assembly101       ✓       ✗       ✓       4,321
      Cross-task       ✗       ✓       ✓       4,713
      RareAct       ✗       ✓       ✓       7,607
      LEGO-VLM (Ours)       ✓       ✓       ✓       35,612
Table 4: Compariable datasets for assembly tasks towards AR training regarding the cabablities of T1, T2, T3. The symbols ✓and ✗indicate the presence or absence of each capability.
COIN provides a strong foundation for sequential task analysis with comprehensive annotations, aiding research in multimodal learning (Tang et al. 2019). HoloAssist captures egocentric human-AI interactions using mixed-reality headsets, offering valuable real-world insights (Wang et al. 2023b). HowTo100M, with its extensive collection of 136 million video clips and transcribed narrations, is highly effective for text-to-video retrieval but lacks fine-grained object annotations (Miech et al. 2019). TEACh focuses on interactive dialogues in domestic environments, improving dialogue modeling but not prioritizing detailed object detection (Padmakumar et al. 2021).
While some datasets excel in action recognition, they often lack the procedural depth necessary for effective training. Assembly101, with over 4,000 toy assembly videos, does not include step-by-step instructional details (Sener et al. 2022). CrossTask facilitates weakly supervised learning by leveraging narrations and step lists but lacks temporal annotations to clarify action sequences (Zhukov et al. 2019). EPIC-KITCHENS provides extensive annotations on unscripted kitchen activities, yet it does not offer structured procedural guidance (Damen et al. 2018). RareAct presents unique interactions, challenging models to interpret complex actions without explicit instructions (Miech et al. 2020).
Stanescu et al. (2023a) introduce a state-aware prior that significantly improves object detection in assembly tasks like furniture and Lego construction. However, it lacks textual descriptions of object states and does not address fine-grained detection (Stanescu et al. 2023b). Meanwhile, MIMIC-IT contributes 2.8 million multimodal instruction-response pairs, enriching conversational modeling but lacking scenario-specific descriptions that would enhance task guidance (Li et al. 2023a).
To bridge these gaps, we conduct a comparative analysis evaluating dataset coverage of essential components—objects, states, and scenes—critical for developing AR-guided assistants. Additionally, we introduce a new simulated dataset featuring nearly 400 objects, systematically alternating between object detection and assembly functions. This dataset is designed to optimize the training and evaluation of advanced VLMs, improving their ability to handle fine-grained assembly tasks.
Towards Fine-grained Vision-Lanuage Models
Recent advancements in vision-language models (VLMs) have broadened their applications, enhancing accessibility, search indexing, and interactive content retrieval (Fan et al. 2024; Zhang et al. 2024; Du et al. 2022; Gan et al. 2022). These models generate captions, describe images (Zhou et al. 2020; Hu et al. 2022), facilitate visual QA (Bazi et al. 2023). For instance, they can accurately identify objects in images based on phrases like “the red car” (Subramanian et al. 2022). Furthermore, these models facilitate conversational interactions about visual elements (Chen et al. 2022), which makes virtual assistants more adept at handling inquiries related to images and videos, which makes virtual assistants more adept at handling inquiries related to images and videos. Additionally, VLMs can support multilingual content (Gwinnup and Duh 2023), benefiting assistive technologies and education (Chi et al. 2020; Wang et al. 2024a).
Despite progress, VLMs often lack fine-grained modeling for scene understanding, object recognition, and error detection. While CLIP demonstrates strong generalization (Radford et al. 2021), models like EfficientVLM (Wang et al. 2023a), MiniGPT-v2 (Chen et al. 2023), and Qwen-VL (Bai et al. 2023) struggle with intricate tasks. OSCAR (Li et al. 2020) and VisionLLM (Wang et al. 2024b) improve image-text alignment but face challenges in context-specific adaptation. The Otter model advances sequential task management yet falls short in detailed analysis (Li et al. 2023b).
Efforts in procedural video representation (Zhong et al. 2023) and instructional task graphs (Ashutosh et al. 2024) highlight potential VLM integration but reveal gaps in error correction. To address these challenges, we propose a vision-language architecture optimized for fine-grained tasks, advancing training and benchmarking of VLMs.
Discussion
Implications and limitations. This work addresses a critical gap in the development of intelligent, instruction-following multimodal assistants by focusing on fine-grained procedural understanding—an ability crucial for equitable access to education and skill development in a technologically evolving society. While our benchmark is partially synthetic and does not yet support 3D scene understanding, it represents an important step toward scalable, accessible AI systems that can interpret complex visual-textual instructions. The implications span beyond research: such systems could democratize hands-on learning by enabling visually impaired users to access assembly tasks through multimodal feedback, and could revolutionize technical education and vocational training by offering consistent, language-grounded support. Our findings further highlight the current limitations of VLMs in precise reasoning and verification, pointing toward future directions in responsible AI development that benefits society at large—including learners, workers, and underserved communities.
Ethical Considerations. We recognize the ethical implications of developing VLMs for user interactions, so addressing these concerns is essential. To ensure ethical standards, we use open-source VLMs as benchmarks and prioritize user-centric design such as inclusiveness for diverse users.
Conclusion
Vision-language models (VLMs) have significantly advanced in recent years, enabling AI systems to interpret and generate textual descriptions of visual content. However, despite their success in general vision-language tasks, these models often struggle with fine-grained understanding, particularly in structured instructional scenarios. This research explores fine-grained vision-language modeling for manual-guided LEGO assembly tasks, focusing on scene understanding, object detection, and state detection—areas where VLMs often struggle. We developed a specialized dataset from LEGO instruction manuals, designing fine-grained tasks to evaluate VLM performance in tracking assembly sequences and interpreting instructions. Through a one-for-all architecture, we assessed existing models and found them lacking in precision for instructional tasks, highlighting the need for improvement. Future work will analyze model failures and develop enhanced VLMs with better accuracy and contextual understanding. Beyond technical advancements, our research aims to empower blind and visually impaired individuals by enabling AI-driven learning tools, promoting greater accessibility and independence.